Assume the random variable x is normally distributed with mean μ=50 and standard deviation σ=7.

Find the indicated probability.

P (x>38)
Answer #1

Solution :

Given that ,

mean = = 50

standard deviation = = 7

P(x >38 ) = 1 - P(x< 38)

= 1 - P[(x -) / < (38-50) /7 ]

= 1 - P(z < -1.71)

Using z table

= 1 - 0.0436

= 0.9564

probability= 0.9564